### v2

# MuSACo: Multimodal Subject-Specific Selection and Adaptation for Expression Recognition with Co-Training

Abstract
Personalized expression recognition (ER) involves adapting a machine learning model to subject-specific data for improved recognition of expressions with considerable interpersonal variability. Subject-specific ER can benefit significantly from multi-source domain adaptation (MSDA) methods – where each domain corresponds to a specific subject – to improve model accuracy and robustness. Despite promising results, state-of-the-art MSDA approaches often overlook multimodal information or blend sources into a single domain, limiting subject diversity and failing to explicitly capture unique subject-specific characteristics. To address these limitations, we introduce MuSACo, a multimodal subject-specific selection and adaptation method for ER based on co-training. It leverages complementary information across multiple modalities and multiple source domains for subject-specific adaptation. This makes MuSACo particularly relevant for affective computing applications in digital health, such as patient-specific assessment for stress or pain, where subject-level nuances are crucial. MuSACo selects source subjects relevant to the target and generates pseudo-labels using the dominant modality for class-aware learning, in conjunction with a class-agnostic loss to learn from less confident target samples. Finally, source features from each modality are aligned, while only confident target features are combined. Experimental results on challenging multimodal ER datasets – BioVid, StressID, and BAH – show that MuSACo outperforms UDA (blending) and state-of-the-art MSDA methods. Our code is available: https://github.com/osamazeeshan/MuSACo
1 Introduction
Figure 1: Comparison of MuSACo against unimodal MSDA and multimodal UDA (blending) ER methods for subject-based adaptation. (a) Unimodal MSDA aligns multiple source subjects to the target within a single modality, which reduces its accuracy. (b) Multimodal UDA with blending incorporates single or blended source domains for target adaptation but does not fully exploit multiple subject-specific diversity. (c) MuSACo selects relevant sources per modality using co-training, and aligns them to the target using both class-aware and class-agnostic losses before fusing modalities for final prediction.
Expression recognition (ER) has achieved significant success in computer vision due to its relevance in numerous real-world applications, such as pain estimation, stress monitoring, and affective computing [aslam2023privileged, aslam2024distilling, li2018deep, waligora2024joint, sharafi2025disentangled]. These models can achieve a high level of performance under supervised settings or when provided with limited labeled data [praveen2020deep, rajasekhar2021deep]. However, their effectiveness mainly depends on the availability of annotated datasets, which are difficult to obtain, and often fail to generalize for subtle expressions and across diverse individuals due to variations in cultural, environmental, and individual expressiveness [jack2009cultural]. Unsupervised domain adaptation (UDA) [gretton2012kernel, long2016unsupervised, xu2018deep, guichemerre2024source] has emerged as a promising direction to address this challenge by adapting a model from a labeled source domain to an unlabeled target domain. In the context of ER, this facilitates personalization, where a model trained on labeled expression data from multiple source subjects is adapted to a new, unlabeled target subject. While existing UDA approaches [han2020personalized, li2018deep, zhu2016discriminative] adapt to the unlabeled target dataset treated as a domain by either using single or blended labeled source subjects, recent advances in multi-source domain adaptation (MSDA) [zeeshan2024subject, zeeshan2025progressive] leverages information from multiple source subjects (domains). This encourages diversity in the source model for adaptation to an unlabeled target subject (domain). This method facilitates more precise and targeted adaptation strategies for developing a personalized ER model (see Fig. 1(a)).
Despite these advances, the majority of subject-specific ER approaches are predominantly constrained to a single modality [han2020personalized, zeeshan2025progressive]. However, human expressions and fine-grained affective cues are often difficult to capture reliably from a single modality alone [liu2021comparing, zhang2020emotion]. To address this limitation, we focus on a multimodal setting that leverages the complementary strengths of multimodal data. Multimodal ER (MER) has been widely studied in the literature [aslam2024distilling, aslam2023privileged, praveen2022joint, tzirakis2017end], typically outperforming unimodal systems by leveraging complementary information from multiple modalities. MER approaches integrate diversity from various modalities, such as visual, textual, audio, or physiological signals [dzedzickis2020human] to enhance model performance. In subject-based ER, most MER methods focus on fully or weakly labeled data [tzirakis2017end, praveen2022joint], hindering the ability of the model to have a better adaptation to unlabeled data. Nevertheless, there are some studies [jimenez2024cfda, wang2022multi] that introduced multiple modalities, specifically for EEG. Eye movement signals with a single source or combine multiple sources into a single domain (see Fig. 1(b)). Their reliance on a single source often fails to provide diversity in source domains. Furthermore, blending multiple sources into a single domain [jimenez2024mmda] cannot explicitly capture the unique subject-specific characteristics of each source subject, which is crucial for subject-specific target adaptation. In this paper, we propose MuSACo, a novel approach for Multimodal Subject-Specific Selection and Adaptation for ER using Co-Training in an MSDA setting (Fig. 1(c)). Our method exploits the complementary information from multiple modalities while selecting the most relevant source subjects for the target. It takes advantage of the unique characteristics of each subject for better performance on the target subject (domain [^1] The terms domain and subject are used interchangeably.). There are two key challenges when considering multiple modalities for subject-specific MSDA: (1) effectively leveraging multiple source subject information within multimodal systems, and (2) achieving cross-domain and cross-modal alignment.
To address the first challenge, we propose the selection of source subjects most relevant to the target subject. While prior work [zeeshan2024subject] has shown that selecting source subjects based on their relevance to the target improves adaptation, it relies solely on a single modality to guide the source selection process. In this paper, we propose a co-training-based strategy for selecting source subjects that leverage multiple modalities, as different modalities complement each other by providing unique characteristics from various perspectives. Thus, selecting sources using co-training benefits the adaptation process. Furthermore, to improve generalization across individuals, we encourage the model to focus on expression-specific features by disentangling identity-related information during source training.
For the second challenge, we first focus on cross-domain alignment. State-of-the-art MSDA methods [kang2020contrastive, scalbert2021multi] highlight the effectiveness of class-aware alignment in bridging dissimilar distributions. This typically involves generating pseudo-labels (PLs) for the target subject. In our proposed method, we extend this by leveraging co-training for the generation of target PLs. Specifically, PLs for unlabeled target data are generated by selecting predictions from the dominant modality (e.g., visual or physiological), chosen by their probability score and threshold. This ensures diversity in modality-specific feature representations and captures complementary aspects of the data. These PLs then guide class-aware alignment between source and target subjects, minimizing distribution mismatch. Although PLs allow improving the model performance, they heavily depend on the choice of a threshold to filter out unreliable samples. In particular, high threshold values ensure the sample quality. While setting a high threshold value often results in discarding potentially useful but less confident samples. To mitigate this, we introduce a class-agnostic loss that aligns non-confident target samples with the source. This encourages the unreliable samples to contribute to the learning process to improve the robustness of the model. For modality alignment, including a fusion module [praveen2022joint, aslam2023privileged] has proven to enhance MER performance. In MuSACo, a fusion module is added to concatenate features from different modalities for each selected source subject. For the target, only reliable samples (selected via co-training) are combined.
The contributions of this paper are summarized as follows. (1) MuSACo: a novel multimodal MSDA method for accurate subject-specific adaptation. It selects relevant source subjects from the target using co-training, leveraging complementary information across modalities. (2) An effective alignment process that relies on co-training to generate confident target PLs for class-aware loss, combined with class-agnostic loss to handle non-confident target samples. (3) Extensive experiments performed on challenging MER datasets (BioVid, StressID, and BAH) show the efficacy of MuSACo. It can outperform the UDA (blending) and state-of-the-art MSDA methods on this multimodal data.
2 Related Work
(a) Multi-Source Domain Adaptation: MSDA techniques for image classification can be classified into discrepancy-based methods [kang2020contrastive, peng2019moment] that reduce the domain shift by estimating the discrepancy between domains. Contrastive-based approaches [kang2020contrastive, scalbert2021multi] generate target pseudo-labels using class-wise clustering, while adversarial and self-supervised methods [zhao2021madan, nguyen2021stem, venkat2020your, deng2022robust] train multiple source-specific classifiers and derive target pseudo-labels through classifier agreement for domain adaptation. Recently, zeeshan2024subject proposed a subject-based MSDA method that treats each subject as a domain to handle many sources. However, existing MSDA approaches remain unimodal (visual-only), motivating our introduction of multimodality for subject-based ER.
(b) Multimodal Expression Recognition: Multimodal ER has been widely studied. Cross-attention fusion has been proposed to combine multiple modalities for expression recognition [praveen2022joint], while LUPI-based methods [aslam2023privileged] and optimal transport distillation [aslam2024distilling] enable learning from multiple modalities but inference with fewer. In domain adaptation, multimodal approaches have been applied to EEG and eye-tracking, e.g., variational autoencoders with cycle consistency [wang2022multi] or CORAL-based alignment with intra/inter-domain losses [jimenez2024cfda]. Building on these, our work leverages visual and physiological modalities in a multi-source DA setting to adapt to unlabeled target subjects.
(c) Disentanglement in Expression Recognition: Disentanglement methods suppress identity or pose factors that hinder ER. GAN-based models such as [jiang2022disentangling, xie2020facial] rely on image reconstruction, making them effective for visual data but unsuitable for non-visual modalities like physiology. In contrast, KNIFE [pichler2022differential] disentangles by minimizing entropy between estimators, a modality-agnostic approach. Motivated by this, MuSACo adopts a similar formulation to decouple identity from expression representations.
(d) Co-training for Domain Adaptation: Co-training is a well-established semi-supervised and unsupervised learning approach [blum1998combining] that is used in many applications, such as semantic segmentation [xia2020uncertainty, gomez2023co], task decomposition [yang2021deep], or in image classification [chen2011co]. It typically trains the model with modalities to produce more reliable pseudo-labels (PLs), while domain adaptation usually trains the model with a single modality that incorporates multiple classifiers to account for different views. Chen et al. [chen2011co] propose a method that introduces two different classifiers trained with different features controlled via a weighting mechanism while generating target pseudo-labels from the most confident classifier. MuSACo introduces a co-training method for selecting source subjects from multimodal data. In addition, we select target confident PLs based on multiple classifiers trained on different modalities.
3 Proposed MuSACo Approach
Figure 2: An overview of MuSACo shown in the particular case of $M$ = 2 modalities (with visual + physiological). First, the similarity between source and target subjects is estimated using the Selection of Source Subjects module, followed by selecting the modality that gives the maximum probability score based on the threshold. Then, Domain Alignment is achieved by the Generation of Target PLs module through co-training. It calculates class-aware loss for each modality, combined with class-agnostic loss to learn from the non-confident target samples. Finally, the Fusion Module is introduced for the Modality Alignment through feature concatenation.
MuSACo seeks to leverage complementary information from multiple modalities within MSDA using co-training to improve target personalization. This is accomplished through a two-step training process. Firstly, the non-discriminative identity information is disentangled from the source subjects while training task-specific backbones using source supervision. This ensures that only discriminative features are retained for recognizing the expression. Second, a co-training strategy is employed to select the most relevant source subjects for the target. For each source-target pair, similarity scores are computed from each modality. Sources that produce high similarity scores in either modality are selected, and a threshold $(\tau_{ss})$ is applied to filter out less relevant sources. Exploiting the trained backbones, we generate target PLs through co-training. The modality with the highest confidence in its prediction is chosen, and the samples are considered reliable if their confidence exceeds a threshold $(\tau_{pl})$. For domain alignment, class-aware alignment is performed on the target samples selected using $\tau_{pl}$ and class-agnostic alignment using non-confident samples. Finally, for modality alignment, features are concatenated from different modalities for every selected source and target subject, which are used to train the fusion module. MuSACo is illustrated in Fig.5.
3.1 Model Architecture
A multimodal multi-source expression classification task is considered, where a model is trained using several labeled source subjects $\mathcal{S}_{m}={\{\mathbf{S}_{m_{1}},\ldots,\mathbf{S}_{m_{a}},\ldots,\mathbf{S}_{m_{D}}\}}$ and a single unlabeled target subject $\mathbf{T}_{m}$, where $m$ is the number of modalities and $a={\{1,\dots,D\}}$ is the number of source subjects. In MuSACo, without loss of generality, we consider two modalities: visual (v) and physiological (p), where $m$ $\in{\{\text{v},\text{p}\}}$. Each source $\mathbf{S}_{m_{a}}$ contains $N^{\textbf{s}}$ labeled samples $\{(\mathbf{x}_{m_{j}}^{\textbf{s}},y_{m_{j}}^{\textbf{s}},\acute{y}_{m_{j}}^{\textbf{s}})|1<j<{N^{\textbf{s}}}$}\}, where $y_{m_{j}}^{\textbf{s}}$ represents expression labels and $\acute{y}_{m_{j}}^{\textbf{s}}$ represents identity labels. We also define expression classes as $c=\{1,\dots,C\}$} The unlabeled target consists of $N^{\textbf{t}}$ samples $\{\mathbf{x}_{m_{j}}^{\textbf{t}}|1<j<{N^{\textbf{t}}}\}$}. Backbones are defined as $\mathcal{B}={\{B_{1},\dots,B_{M}\}}$, where each $B$ belongs to its respective modality. Accordingly, we consider two backbones that correspond to $B_{\text{v}}$ (visual) and $B_{\text{p}}$ (physiological) modalities. It takes input from $\mathbf{x}_{m}^{\textbf{s}}$ and $\mathbf{x}_{m}^{\textbf{t}}$ and return embeddings of source $\mathbf{h}^{\textbf{s}}_{m}=\mathbf{h}_{\text{v}}^{\textbf{s}},\mathbf{h}_{\text{p}}^{\mathbf{s}}$ and target $\mathbf{h}^{\textbf{t}}_{m}=\mathbf{h}_{\text{v}}^{\textbf{t}},\mathbf{h}_{\text{p}}^{\textbf{t}}$. Finally, the fusion module $F$ is a two-layer perceptron that takes input extracted from $\mathcal{S}_{m}$ and $\mathbf{T}_{m}$ to perform feature fusion for the alignment of modalities.
3.2 Training Backbones with Disentanglement
Two backbones are trained for each modality – visual $B_{\text{v}}$ and physiological $B_{\text{p}}$ – using multiple labeled source subjects $\mathcal{S}_{m}$. To learn expression-discriminative representations, each modality is supervised using a cross-entropy:
$\mathcal{L}_{m}^{\textbf{s}}=\frac{1}{D}\frac{1}{N^{\textbf{s}}}\sum_{i=1}^{D}\sum_{j=1}^{N^{\textbf{s}}}\text{CE}(\mathbf{y}_{m_{j}}^{i},f_{m}(\mathbf{h}_{m_{j}}^{i}))$ (1)
where $\mathbf{h}_{m_{j}}^{i}$ is the embedding from modality $m$, $\mathbf{y}_{m_{j}}^{i}$ is the ground-truth expression label, and $f_{m}$ is the modality-specific expression classifier. The loss $\mathcal{L}_{m}^{\textbf{s}}$ is computed independently for each modality to ensure effective learning of modality-specific features.
Disentangled Identity-Information. Our approach is inspired by KNIFE [pichler2022differential], a differentiable entropy estimator, to disentangle identity information from expression-relevant features in each modality. We estimate the marginal entropy $H(\mathbf{h}_{m}^{\mathbf{s}})$ and the conditional entropy $H(\mathbf{h}_{m_{j}}^{\mathbf{s}}|\acute{Y}_{m_{j}}^{\mathbf{s}})$. For each sample in the source subjects, the identity labels $\acute{y}_{m_{j}}^{\textbf{s}}$ are converted into one-hot representations $\acute{Y}_{m_{j}}^{\textbf{s}}\in R^{k}$ and define the disentanglement loss as:
$\mathcal{L}_{m}^{d}=H(\mathbf{h}_{m}^{\textbf{s}})+H(\mathbf{h}_{m_{j}}^{\textbf{s}}|\acute{Y}_{m_{j}}^{\textbf{s}}).$ (2)
The classification loss $\mathcal{L}_{m}^{s}$ and disentanglement loss $\mathcal{L}_{m}^{d}$ are jointly optimized to preserve task-relevant features while removing identity information [^2] More details are provided in Section 2 of the supplementary material.. Importantly, since $\mathcal{L}_{m}^{d}$ is entropy-based and modality-agnostic, it is applied independently to each modality (visual, physiological, and audio), ensuring that subject-specific identity cues do not interfere with expression-related representations and improving robustness across modalities.
3.3 Domain Adaptation via Co-training
(a) Selection of Source Subjects. Selecting the most relevant source subjects from $\mathcal{S}_{m}$ is a critical step in MuSACo method, where the model is trained on diverse sources to ensure robustness. However, when adapting to a target subject $\mathbf{T}_{m}$ captured in a specific environment and cultural context (e.g., an adult Caucasian male), the source subjects must align with these characteristics to ensure effective adaptation. To address this, we leverage our multimodal system to define selection criteria based on co-training, which ensures that the chosen sources are highly relevant to the target subject. Given two modalities visualvisual and physiologicalphysiological, estimate the cosine similarity between $\mathcal{S}$ and $\mathbf{T}$ using embeddings $\mathbf{h}_{m}^{\textbf{s}}$ and $\mathbf{h}_{m}^{\textbf{t}}$,
$\mathbf{Q}_{m}(\mathcal{S},\mathbf{T})=\frac{\mathbf{h}_{m}^{\textbf{s}}\cdot\mathbf{h}_{m}^{\textbf{t}}}{\|\mathbf{h}_{m}^{\textbf{s}}\|_{2}\cdot\|\mathbf{h}_{m}^{\textbf{t}}\|_{2}}$ (3)
where $||.||$ is the euclidean norm calculated for $\mathbf{h}_{m}^{\textbf{s}}$ and $\mathbf{h}_{m}^{\textbf{t}}$. $\mathbf{Q}_{m}(\mathcal{S},\mathbf{T})$ is estimated for each source and target pairs as,
$\begin{split}P_{m}=[\mathbf{Q}_{m}^{\text{c}}(\mathbf{S}_{m_{1}},\mathbf{T}_{m}),\ldots,\mathbf{Q}_{v}^{\text{c}}(\mathbf{S}_{m_{D}},\mathbf{T}_{m})]\end{split}$ (4)
Here, $P_{m}$ represents a dictionary list containing all pairwise distances computed using cosine similarity for each modality. Due to the distinct architectures of the modalities, such as a 2D CNN for the visual modality and a 1D CNN for the physiological modality, we normalize $P_{m}$ to the range $[0,1]$ to ensure comparability across modalities. We define $P_{\text{v}}$ and $P_{\text{p}}$ as the similarity scores for the visual and physiological modalities, respectively. These scores are merged into a combined set $\widehat{P}=(P_{\text{v}}\bigcup P_{\text{p}})$, from which we select source subjects whose distances exceed a predefined threshold $\tau_{ss}$. The selection process is formalized as:
$\widetilde{\mathcal{S}}_{m}=\{\mathcal{S}_{m_{i}}:\widehat{P}_{i}>\tau_{ss}\}\quad\forall i\in\left\{1,\dots,D\right\}$ (5)
The selected source subjects for each modality form a subset $\widetilde{\mathcal{S}}_{m}$, where $\widetilde{\mathcal{S}}_{m}\subseteq\mathcal{S}_{m}$. This selection process ensures that only the most relevant source subjects are used for the target adaptation.
(b) Generation of Target Pseudo-Labels (PLs). In MSDA, reliable PLs determine the performance of a model on an unseen target subject [zeeshan2024subject]. MuSACo benefits from the multimodal by selecting confident target PLs using co-training. Given two backbones, $B_{\text{v}}$ and $B_{\text{p}}$ We introduce samples $\mathbf{x}_{m}^{\textbf{t}}$ from $\mathbf{T}_{m}$ to produce softmax probabilities with respect to each modality, $p_{\text{v}}=\sigma(\mathbf{h}_{\text{v}}^{\textbf{s}})$ and $p_{\text{p}}=\sigma(\mathbf{h}_{\text{p}}^{\textbf{s}})$, taking the maximum probability $\widehat{p}=max(p_{\text{v}},p_{\text{p}})$ and applying a predetermined threshold $\tau_{pl}$ to select the confident samples, defined as:
${\textbf{T}}_{m}^{l}=\left\{({\hat{x}_{m_{j}},\widehat{y}_{m_{j}}})|{\widehat{p}>\tau_{pl}},{j\in{\{1\leq j\leq N^{\textbf{t}}\}}}\right\}$ (6)
where $\textbf{T}^{l}_{m}$ represents the set of confident target samples for $m$ modalities, we also store the confident classes $C={\{{y_{i}|(x_{i},y_{i})\in\textbf{T}}_{m}^{l}\}}$. To differentiate between confident and non-confident samples, we create ${\textbf{T}}^{u}_{m}$, defined as:
${\textbf{T}}_{m}^{u}=\left\{({{x}_{m_{j}}})|{\widehat{p}\leq\tau_{pl}},{j\in{\{1\leq j\leq N^{\textbf{t}}\}}}\right\}$ (7)
where $\textbf{T}_{m}^{u}$, defined the non-confident target samples, for the estimation of class-agnostic alignment (Section 3.3).
(c) Class-Aware Domain Alignment. We exploit class information to perform intra-class and inter-class alignment across domains. For intra-class alignment, samples from the same class are encouraged to reside in the same distribution space. Conversely, for inter-class alignment, samples from different classes are pushed further apart in the feature space. Class-aware alignment is computed using maximum mean discrepancy (MMD) [sejdinovic2013equivalence], a statistical measure for estimating the disparity between two distributions. To sample out examples from the same class, we perform class-aware sampling on source and target subjects. The classes that belong to $\mathbf{T}^{l}_{m}$ were denoted as $\widetilde{C}$. We sample out examples from every subject belonging to multiple labeled sources. In each mini-batch, examples from every source and target are extracted for the same class. The intra-class discrepancy is defined as:
$\begin{split}D_{\text{ita}}(\widetilde{\mathcal{S}}_{m},\mathbf{T}^{l}_{m})=\frac{1}{D}\frac{1}{\widetilde{C}}\sum_{a=1}^{D}\sum_{{c}=1}^{\widetilde{C}}\text{MMD}(\mathbf{h_{m_{a}}^{c}},\mathbf{h_{m}^{c}})\end{split}$ (8)
where, $\mathbf{h_{m_{a}}^{c}}$ extracts features from each source subject, $\mathbf{h_{m}^{c}}$ are the features extracted from the target subject, and $\widetilde{C}$ are the confident expression classes. Inter-class is defined as:
$\begin{split}D_{\text{itr}}(\widetilde{\mathcal{S}}_{m},\mathbf{T}^{l}_{m})=\frac{1}{D}\frac{1}{\widetilde{C}(\widetilde{C}-1)}\sum_{a=1}^{D}\sum_{c=1}^{\widetilde{C}}\sum_{\binom{\acute{c}=1}{\acute{c}\neq c}}^{\widetilde{C}}\text{MMD}(\mathbf{h_{m_{a}}^{c}},\mathbf{h}_{m}^{\acute{c}})\end{split}$ (9)
where MMD is computed from the opposite classes $\acute{c}$ for every source with the target. Alignment loss is estimated for every modality (visual and physio), ensuring that source and target align in their respective distribution space.
(d) Class-Agnostic Domain Alignment. Most MSDA methods [kang2020contrastive, scalbert2021multi] employ class-aware alignment to bridge dissimilar distributions, often relying on PLs for the target subject. However, these pseudo-labels are prone to noise, and while denoising techniques have been proposed [scalbert2021multi, xu2018deep, deng2022robust, ren2022multi], they typically use pre-determined thresholds to retain only the most confident pseudo-labels (CPL). This reliance on thresholds can limit performance, especially in ER, where subtle expressions may lead to discarding useful but less confident samples. To address this, a domain-agnostic loss is introduced that aligns non-confident samples to ensure that all target samples contribute to the adaptation process. This approach maximizes the utilization of available data for improving the target adaptation. The domain-agnostic loss is defined as:
$D_{\text{agn}}(\widetilde{\mathcal{S}}_{m},\mathbf{T}^{u}_{m})=\frac{1}{D}\sum_{a=1}^{D}\text{MMD}(\mathbf{h}_{m_{a}}^{\textbf{s}},\mathbf{h}_{m}^{\textbf{t}})$ (10)
where $\mathbf{h}_{m_{a}}^{s}$ is the feature vector extracted from $\widetilde{\mathcal{S}}_{m}$ subjects that contain samples not belonging to $\widetilde{C}$. The total alignment loss for both modalities is calculated as:
$\begin{split}\mathcal{L}_{\text{agn}}=D_{\text{agn}}(\widetilde{\mathcal{S}}_{\text{v}},\mathbf{T}^{u}_{\text{v}})+D_{\text{agn}}(\widetilde{\mathcal{S}}_{\text{p}},\mathbf{T}^{u}_{\text{p}})\end{split}$ (11)
$\begin{split}\mathcal{L}_{\text{aw}}=[D_{\text{ita}}(\widetilde{\mathcal{S}}_{\text{v}},\mathbf{T}^{l}_{\text{v}})-D_{\text{itr}}(\widetilde{\mathcal{S}}_{\text{v}},\mathbf{T}^{l}_{\text{v}})]+\\ [D_{\text{ita}}(\widetilde{\mathcal{S}}_{\text{p}},\mathbf{T}^{l}_{\text{p}})-D_{\text{itr}}(\widetilde{\mathcal{S}}_{\text{p}},\mathbf{T}^{l}_{\text{p}})]\end{split}$ (12)
where $\mathcal{L}_{\text{agn}}$ and $\mathcal{L}_{\text{aw}}$ are class-agnostic and class-aware losses respectively. For each modality, the intra-class and class-agnostic discrepancy is minimized, while the inter-class discrepancy is maximized.
(e) Modality Alignment. In multimodal settings, aligning modalities by adding a fusion network has proven to be very successful in enhancing model performance [praveen2022joint, aslam2023privileged]. In MuSACo method, to align the modalities, a fusion module $F$ is introduced that projects the combined modality embeddings into a shared feature space. To train $F$ for multiple source subjects, we use the embeddings $\mathbf{h}_{\text{v}}^{\textbf{s}}$ and $\mathbf{h}_{\text{p}}^{\textbf{s}}$ to perform feature concatenation $\mathbf{h}^{\textbf{s}}=\mathbf{h}_{\text{v}}^{\textbf{s}}\oplus\mathbf{h}_{\text{p}}^{\textbf{s}}$, and pass through $F$ to calculate a cross-entropy loss:
$\mathcal{L}_{\text{}}^{\textbf{s}}=\frac{1}{D}\frac{1}{N^{\textbf{s}}}\sum_{i=1}^{D}\sum_{j=1}^{N^{\textbf{s}}}\text{CE}(\mathbf{y}_{j}^{i},F(\mathbf{h}_{j}^{i}))$ (13)
where $D$ is the source subject that contains the $N^{\textbf{s}}$ number of samples. Modality alignment is estimated for each $\widetilde{\mathcal{S}}$ and $\mathbf{T}$ subject individually. We only calculated for the confident samples for the target $\mathbf{T}^{l}_{m}$, the embeddings are extracted $\mathbf{h}_{\text{v}}^{\textbf{t}}$ and $\mathbf{h}_{\text{p}}^{\textbf{t}}$, and applied concatenation function $\mathbf{h}^{\textbf{t}}=\mathbf{h}_{\text{v}}^{\textbf{t}}\oplus\mathbf{h}_{\text{p}}^{\textbf{t}}$, then $F$ is applied to calculate the cross-entropy loss:
$\mathcal{L}_{\text{unsup}}^{\textbf{t}}=\frac{1}{N_{l}^{\textbf{t}}}\sum_{j=1}^{N_{l}^{\textbf{t}}}\text{CE}(\widehat{\mathbf{y}}_{j},F(\mathbf{h}_{j}^{\textbf{t}}))$ (14)
where $N_{l}^{\textbf{t}}$ is the total number of reliable samples, and $\widehat{\mathbf{y}}_{j}$ is the CPL. The final objective is to jointly optimize all four losses $\mathcal{L}=\mathcal{L}^{\textbf{s}}+\gamma\mathcal{L}_{\text{unsup}}^{\textbf{t}}+\alpha\mathcal{L}_{\text{agn}}+\beta\mathcal{L}_{\text{aw}}$, where $\gamma$, $\alpha$, and $\beta$ hyper-parameters weight the contribution of each loss.
4 Results and Discussion
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg Acc Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
37.1
41.3
39.1
27.2
33.2
35.2
26.3
40.3
28.6
29.6
34.2
25.6
28.4
33.4
39.2
28.6
31.2
36.6
35.6
37.3
44.5
22.1
33.9
27.8
33.6
35.6
28.8
19.7
25.2
20.1
28.8
34.6
32.5
20.8
26.4
28.1
$M$-UDA
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
39.4
32.6
34.1
43.5
28.4
24.1
26.3
29.9
31.5
29.2
31.5
37.6
33.4
32.1
36.7
38.2
37.6
29.6
34.2
40.0
31.5
26.8
33.5
22.3
34.2
28.4
43.4
38.6
26.6
23.6
34.1
40.0
34.3
23.4
28.6
47.6
27.4
22.6
31.2
25.8
32.4
27.2
33.4
36.3
27.4
25.3
28.2
28.5
$M$-MSDA
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
CMSDA [scalbert2021multi]
MuSACo (MSDA)
43.2
43.5
39.4
49.2
32.2
42.2
39.4
43.3
27.2
28.5
29.4
47.3
33.6
22.5
36.6
38.2
33.4
29.1
39.2
40.1
32.5
32.5
39.1
44.3
36.6
40.6
38.4
46.4
32.4
41.2
35.7
45.2
41.5
38.0
45.2
50.0
33.8
28.4
25.6
34.3
34.6
34.7
36.8
43.8
27.3
28.4
30.0
35.7
Upper Bound Fine-tuning 83.6 78.5 70.0 71.6 71.5 64.5 76.8 76.9 78.5 64.8 73.6 69.1
Table 1: Accuracy and Average F1 score of MuSACo with baselines against multimodal UDA and MSDA methods on BioVid. Best Avg Acc and Avg F1 per section are highlighted in bold (ignoring supervised fine-tuning).
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg Acc Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
71.1
57.7
68.3
23.2
81.1
30.4
51.4
60.0
54.2
63.3
28.2
57.6
55.6
60.0
65.3
63.2
69.7
67.8
72.7
84.2
79.7
48.7
51.1
48.6
60.0
34.6
61.2
59.6
29.7
59.4
56.9
55.6
59.3
47.9
34.8
50.5
$M$-UDA
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
65.3
58.1
66.3
71.2
65.3
68.1
52.3
53.4
73.3
72.6
71.6
69.7
66.4
59.9
74.5
70.0
67.3
67.9
71.9
73.5
72.3
70.7
74.5
68.7
73.2
76.4
81.2
82.4
55.3
59.5
57.2
60.0
70.4
60.3
68.4
72.7
57.2
49.4
57.2
61.2
66.6
64.2
67.5
68.3
52.9
52.5
53.0
60.7
$M$-MSDA
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
69.3
70.8
71.2
81.2
74.9
81.2
67.2
73.6
73.6
59.3
70.6
74.6
62.3
59.8
73.5
74.2
77.9
78.5
83.3
71.8
86.3
60.0
49.9
60.0
66.8
63.9
79.2
59.3
69.8
71.5
68.3
68.3
74.9
53.2
54.1
62.2
Upper Bound Fine-tuning 83.4 84.5 98.2 98.1 89.7 87.6 90.0 89.6 89.4 95.7 90.6 80.4
Table 2: Accuracy and Average F1 score of MuSACo with baselines, multimodal UDA and MSDA methods on StressID adapted to 10 unlabeled target subjects. Best Avg Acc and Avg F1 per section are highlighted in bold.
Methods Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg Acc Avg F1
Visual-only 33.3 86.0 47.1 70.5 73.0 62.0 38.2
Audio-only 35.3 35.4 44.1 65.4 60.3 48.1 36.1
Fusion 36.8 81.7 47.6 73.6 77.6 63.5 38.4
MMD [sejdinovic2013equivalence] 29.9 81.0 53.4 68.8 77.6 62.1 38.8
MuSACo (UDA) 35.4 86.6 46.7 73.5 79.4 64.3 38.7
Sub-basedtop-k [zeeshan2024subject] 33.4 84.5 54.2 71.4 79.5 64.6 38.9
MuSACo (MSDA) 39.7 95.0 54.2 79.3 81.2 69.9 41.2
Table 3: Performance of MuSACo against baselines, multimodal UDA, and MSDA methods on the BAH dataset.
4.1 Experimental Protocol
Datasets. MuSACo was evaluated on three challenging multimodal ER datasets for subject-based adaptation. BioVid Heat Pain (Part A) [walter2013biovid] contains 87 subjects with five pain levels; following [zeeshan2024subject], we use 77 as sources and 10 as targets. StressID [chaptoukaev2023stressid] comprises 65 participants in controlled settings. For our experimental setup, 44 source and 10 target subjects were used with both visual and physiological (EDA) modalities. Behavioural Ambivalence/Hesitancy (BAH) [gonzalez2025bah] consists of 224 participants from diverse demographics in uncontrolled, real-world settings. Our analysis utilizes 143 of these individuals as source subjects and five as targets with visual and audio modalities.
Implementation Detail. In addition to the visual modality, the physiological modality, specifically electrodermal activity (EDA), is incorporated for both the BioVid and StressID datasets. For visual, ResNet18 [he2016deep] is used, following the same protocol as [zeeshan2024subject]. Additional experiments using an alternative visual backbone are reported in the suppl material. For the physio modality, an LSTM-based 1D-CNN network, and for the audio modality, PANN with CNN14 [kong2020panns] model is used for the BAH dataset. All the target subjects are split into train/val/test sets, and all the results are reported using a target test set. For selection of source subjects, $\tau_{ss}$ is set to 0.55 (Sec. 4.4-b). Following prior work [deng2022robust, scalbert2021multi, ren2022multi], we set the initial $\tau_{pl}$ to 0.95 and gradually decrease it by $0.01$ every $N$ epochs to incorporate more samples as training stabilizes, with this choice validated as well as further datasets and implementation details, provided in the suppl material.
Baseline Methods. The reported results correspond to the top-1% multimodal target accuracy across all target subjects. The evaluation begins with the Lower Bound, where a model is trained solely on labeled source subjects and directly evaluated on the target test set without any target supervision. Results are reported for visual-only, physio-only, or audio-only, and fusion (vision+physio) or (vision+audio) modalities. Next, the MM-UDA (blending) setting blends all the source subjects into a single domain, then adapts to the unlabeled target subject using various UDA techniques, including DANN [ganin2016domain], CDAN [long2018conditional], MMD [sejdinovic2013equivalence], and MuSACo (UDA), which only use co-training to generate target pseudo-labels. Furthermore, MuSACo (MSDA) is compared against MM-MSDA approaches, including CAN [kang2020contrastive], Sub-based MSDA, and CMSDA [scalbert2021multi]. In all methods, we use the same backbones for every modality to ensure a fair comparison with our proposed approach. Finally, results for an Upper Bound are reported, where models are fine-tuned using labeled data from the target subject.
4.2 Comparison with State-of-the-Art Methods
Tab. 1 shows results on the BioVid dataset. All domain adaptation approaches outperform the lower bound (source-only) baselines (visual and both), confirming the effectiveness of multimodal domain adaptation in the context of pain estimation. Notably, the physio-only baseline outperforms many standard UDA and MSDA methods, achieving an average accuracy of 34.5%. This reflects findings in prior work [werner2017analysis, waligora2024joint, aslam2024distilling], where physio signals are shown to be strong indicators of pain intensity, especially in subtle expressions where visual cues are often weak or ambiguous. Similarly, our visual-only and fusion baselines yield weaker accuracy than physio-only. Nonetheless, fusion improves recall on harder classes, resulting in a higher Avg F1 score (28.1%) compared to physio-only (26.4%), despite its slightly lower accuracy. Among all methods, our approach achieves the highest overall performance. MuSACo (UDA) outperforms both the lower-bound and UDA baselines and is competitive with MSDA methods. Further, MuSACo (MSDA) achieves even better results, with a 9.3% gain over physio-only and 7.5% over MuSACo (UDA). It also outperforms MSDA baselines, with an average gain of 9.2% over CAN [kang2020contrastive] and Sub-basedtop-k [zeeshan2024subject], and 7% over CMSDA [scalbert2021multi]. This shows our method effectively exploits modality complementarity, adapting to subject-specific signals in subtle pain recognition.
Figure 3: Left: t-SNE visualizations. The source-only produced indistinguishable feature class clusters. Sub-based MSDA [zeeshan2024subject] reduces some noise by separating classes to some extent. MuSACo creates more separable clusters for each class. Right: Visualization of selected and non-selected source subjects for a reference target subject, with similarity scores from visual and physiological modalities.
Tab. 2 presents the performance on the StressID dataset. As expected, training without adaptation (lower bound) yields significantly lower accuracy across most subjects. Notably, fusing both modalities in the lower bound produces more stable performance compared to using either modality alone. This is due to the complementary nature of the modalities; when the visual modality underperforms, the physio often compensates, resulting in more balanced performance across subjects. Among existing methods, CAN achieves the highest performance on Sub-2, 8, whereas Sub-basedtop-k gains the performance on Sub-3 that is comparable with our method. MuSACo (UDA) improved the performance on 3 out of 10 subjects while, on average, attaining a similar performance as MSDA methods. MuSACo (MSDA) outperformed every other method on all target subjects, with an overall gain of 15.7% from lower-bound (both), 7.4% from MMD, 6.7% from MuSACo (UDA) and CAN, and 6.6% from Sub-based MSDA. Tab. 3 reports results on the BAH dataset. It can be seen that the visual modality is much stronger than audio, but fusion leverages complementary information across both modalities to achieve slight gains in performance over unimodal baselines. Among the UDA methods, MMD improves over the lower bound, while MuSACo (UDA) further stabilizes adaptation across subjects. In the MSDA setting, Sub-based demonstrates competitive results; however, MuSACo (MSDA) consistently outperforms all baselines, surpassing the lower-bound fusion by 6.4% in average accuracy and 2.3% in F1 score. Subject-wise F1 scores were provided in suppl. material.
4.3 Visualizations
In Fig. 3 (left), t-SNE [van2008visualizing] shows how different methods separate class embeddings. The source-only model yields noisy, overlapping features. Subject-based MSDA [zeeshan2024subject] improves clustering but struggles with overlaps (e.g., PA3 in PA4). In contrast, MuSACo better clusters same-class samples and separates opposing ones. Fig. 3 (right) shows the closest and furthest source subjects to a target (Sub-10, BioVid), based on similarity scores from visual and physiological modalities. A source is selected if either modality exceeds the threshold $\tau_{ss}$, e.g., Sub-1/2 (visual) and Sub-3/4 (physio). Sub-5 to Sub-8 are excluded as both scores fall below $\tau_{ss}$ (see Sec. 4.4(b) for threshold selection).
4.4 Ablation Studies
(a) Impact of difference loss components. An incremental ablation study is conducted by gradually introducing loss terms under the same configuration and weighting hyperparameters as the final model. Tab. 4 shows that adding the target loss $\mathcal{L}^{t}_{unsup}$ significantly improves Accuracy and F1 scores over the source-only baseline. Further gains are achieved by incorporating the class-aware loss $\mathcal{L}_{aw}$ and class-agnostic loss $\mathcal{L}_{agn}$, both of which contribute to improving the adaptation process in the absence of target labels. The combination of all loss terms achieves the highest performance, reaching 43.8% Accuracy and 35.7% F1. (b) Impact of selection of source subjects threshold. In Fig. 4, setting $\tau_{ss}$=0 disables the selection mechanism, including all source subjects (77 for Biovid), achieving an average accuracy of 37.6%. As $\tau_{ss}$ increases, fewer but more relevant source subjects are selected based on similarity scores, improving performance. The best accuracy of 39.8% is achieved at $\tau_{ss}$=0.55, where only 21 source subjects are selected on average. This shows the effectiveness of co-training in identifying complementary informative sources for better target adaptation. Subject-wise results are provided in the suppl. material.
Figure 4: Source subject threshold ($\tau_{ss}$) selection based on the accuracy. The red circle highlights the selected $\tau_{ss}=0.55$.
$\mathcal{L}^{s}$ $\mathcal{L}^{t}_{unsup}$ $\mathcal{L}_{aw}$ $\mathcal{L}_{agn}$ Avg Acc Avg F1
32.5 28.1
36.6 29.2
38.1 29.7
39.3 31.2
43.8 35.7
Table 4: Impact of individual loss components. Average performance across all subjects of the BioVid dataset.
5 Conclusion
Domain adaptation methods typically focus on individual modalities, or in the case of multimodal MSDA, they aggregate data from all source subjects, losing critical subject-specific cues necessary for effective adaptation to unlabeled target data. In this paper, MuSACo, a multimodal subject-specific MSDA is introduced that effectively leverages subject-level information for a personalized adaptation. In particular, MuSACo selects the most similar source subjects based on multimodal similarity and applies co-training to generate PLs for the target. Class-aware and class-agnostic losses are combined to leverage both confident and uncertain target samples effectively. Final alignment is achieved via feature concatenation across modalities. MuSACo is validated on challenging MER datasets, showing consistent gains across methods. MuSACo supports personalized modeling by adapting to each target subject through relevant sources, making it well-suited for health-related applications. Future work will explore test-time adaptation to reduce per-subject retraining and enable faster subject-specific adaptation.
Acknowledgment
This research was partially supported by the Natural Sciences and Engineering Research Council of Canada, Fonds de recherche du Québec – Santé, Canada Foundation for Innovation, and the Digital Research Alliance of Canada.
6 Supplementary
7 Additional Implementation Details
For visual, ResNet18 [he2016deep] network, physio network based on an LSTM-based 1D-CNN network, consisting of two convolutional layers, one LSTM layer, and one fully connected layer. For the audio modality, we employ the Pretrained Audio Neural Network (PANNs) [kong2020panns] framework, specifically the CNN14 architecture, pretrained on AudioSet. Additionally, the expression head is constructed using two 2-MLP layers for each network. A batch size of 32 for the target subject with images resized to $\times100$ resolution, and the model is trained for 20 epochs.
7.1 Algorithm
(a) Selection of Subject-Specific Sources. Algorithm 1 shows the subject-specific selection of source subjects with co-training. Given source subjects $\mathcal{S}$ and specific target subject $\mathbf{T}$, extracted features using visual $B_{\text{v}}$ and physiological $B_{\text{p}}$ encoders. Construct a similarity metric $z^{v}$ and $z^{p}$ by measuring CosineSimilarity between every source subject and target. To estimate the maximum score, we normalize the similarities, merge them, and sort them in descending order, followed by a threshold that selects the most relevant subjects from the target.
Algorithm 1 Subject-Specific Selection of Source Subjects
$\mathcal{{S}}$: labeled source subjects, $\mathbf{T}$: unlabeled target domain, $B_{\text{vis}}$: visual encoder, $B_{\text{phy}}$: physio encoder, $\tau_{ss}$: threshold to select relevant source subjects
1:Initialize: $\mathbf{z}^{v}\leftarrow[]$, $\mathbf{z}^{p}\leftarrow[]$
2: Decompose $\mathbf{T}$ into $\mathrm{T}^{v}$ and $\mathrm{T}^{p}$
3:# Extract target features
[…]
5:for each domain $\mathbf{S}_{i}\in\mathcal{S}$ do
6:   Decompose $\mathbf{S}_{i}$ into $\mathrm{S}^{v}$ and $\mathrm{S}^{p}$
7:  # Extract source features
8:  $\mathrm{X}_{s}^{v}\leftarrow{B}_{vis}(\mathrm{S}^{v})$, $\mathrm{X}_{s}^{p}\leftarrow{B}_{phy}(\mathrm{S}^{p})$
9:  # Compute similarities
10:  $\mathrm{z}_{i}^{v}\leftarrow\cos(\mathrm{X}_{s}^{v},\mathrm{X}_{t}^{v})$
11:  $\mathrm{z}_{i}^{p}\leftarrow\cos(\mathrm{X}_{s}^{p},\mathrm{X}_{t}^{p})$
12:  # Append similarities
13:  $\mathbf{z}^{v}\leftarrow\mathbf{z}^{v}\cup\mathrm{z}_{i}^{v}$, $\mathbf{z}^{p}\leftarrow\mathbf{z}^{p}\cup\mathrm{z}_{i}^{p}$
14:end for
[…]
16:Merge similarities: $\mathbf{z}\leftarrow\mathrm{merge}(\mathbf{z}^{v},\mathbf{z}^{p})$
17:Sort $\mathbf{z}$ in descending order
18:Select relevant sources : $\widetilde{\mathcal{S}}\leftarrow\tau_{ss}(\mathbf{z})$
(b) Training Protocol of MuSACo. In algorithm 2, we show our training protocol of adapting to an unlabeled target subject. All the equation numbers referenced here correspond to those in the main paper.
Algorithm 2 Training protocol of MSACo
$\widetilde{\mathcal{{S}}}$: selected labeled source subjects
$\mathbf{T}$: unlabeled target domain
1:for epoch do
2:   Perform co-training on $\mathbf{T}$ to generate pseudo-labels (after every n epochs)
3:  for iteration do
4:   # Class-aware alignment
5:    Class-aware domain sampling of $\widetilde{\mathcal{S}}$ and $\mathbf{T}$
6:   Estimate intra-class discrepancy using (Eq: 8)
7:   Estimate inter-class discrepancy using (Eq: 9)
8:   Compute class-aware loss $\mathcal{L}_{aw}$ (Eq: 12)
9:   # Domain-agnostic alignment
10:   Domain-agnostic sampling of $\widetilde{\mathcal{S}}$ and $\mathbf{T}$
11:   Estimate domain agnostic using (Eq 10)
12:   Compute domain-agnostic loss $\mathcal{L}_{agn}$ (Eq: 11)
13:   # Modality alignment
14:    Perform feat concatenation of source subjects between modalities $\mathbf{h}_{\text{v}}^{\textbf{s}}$ and $\mathbf{h}_{\text{p}}^{\textbf{s}}$
15:    Perform feat concatenation of target subject between modalities $\mathbf{h}_{\text{v}}^{\textbf{t}}$ and $\mathbf{h}_{\text{p}}^{\textbf{t}}$
16:   Compute modality align loss $\mathcal{L}_{\text{}}^{\textbf{s}}$ using (Eq: 13)
17:   Compute modality align $\mathcal{L}_{\text{unsup}}^{\textbf{t}}$ using (Eq: 14)
18:  end for
19:end for
7.2 Hyperparameters
Training of backbones. momentum=0.9, weight decay=5e-4, stochastic gradient descent (SGD) optimizer [ruder2016overview], learning rate=1e-4 with lr scheduler (eta_min=0.00002). Weighting parameters. In MuSACo, we give different weights to different loss functions. For the contribution of class-agnostic loss $\gamma=0.5$, and for class-aware loss $\alpha=0.1$. Disentanglement. Knife is very sensitive to hyperparameters; we have explored several parameters to make it work with the expression recognition task. The most critical parameters that is selected for our experiments are: zd_dim=1024, zc_dim=77, hidden_state=512, layers=3, nb_mixture=10, with learning-rate=0.01.
8 Detail on Training the Source Backbones
Figure 5: Training of disentanglement with knife-loss and identity-head.
Training of the source backbones (visual and physiological) involves disentangling the identity-related features from the expression task. Our method is inspired by KNIFE[pichler2022differential], a fully differentiable entropy estimator, which we adapted for disentanglement in a multi-modal framework. The KNIFE estimator optimizes the backbones (visual and physiological) by decoupling non-task-related information through individual modality-specific gradient-based optimization. We leverage the KNIFE estimator to minimize the mutual information between modalities in eliminating identity features and enhancing the disentanglement of task-relevant features across distributions. We first estimate the marginal entropy of embedding $\mathbf{h}_{m}^{\textbf{s}}$ as,
$H(\mathbf{h}_{m}^{\textbf{s}})=-\mathbb{E}[MM{p}(\mathbf{h}_{m}^{\textbf{s}})]$ (15)
where $\mathbb{E}$ is the expectation over the distribution $\mathbf{h}_{m}^{\textbf{s}}$ and $p(.)$ is the probability density. To estimate the conditional entropy between the features and identities, For each sample in the source subjects, we convert the identity labels $\acute{y}_{m_{j}}^{\textbf{s}}$ into one-hot representations $\acute{Y}_{m_{j}}^{\textbf{s}}\in R^{k}$. Then, conditional entropy is calculated as,
$H(\mathbf{h}_{m_{j}}^{\textbf{s}}|\acute{Y}_{m_{j}}^{s})=-\mathbb{E}[\log{p}(\mathbf{h}_{m_{j}}^{\textbf{s}}|\acute{Y}_{m_{j}}^{\textbf{s}})]$ (16)
where $p(.,.)$ is the conditional probability density between features $\mathbf{h}_{m_{j}}^{\textbf{s}}$ and prediction $\acute{Y}_{m_{j}}^{\textbf{s}}$. The total learning loss is calculated as:
$\mathcal{L}_{m}^{d}=H(\mathbf{h}_{m}^{\textbf{s}})+H(\mathbf{h}_{m_{j}}^{\textbf{s}}|\acute{Y}_{m_{j}}^{\textbf{s}}).$ (17)
By minimizing $\mathcal{L}_{m}^{d}$, the model is encouraged to decouple the information associated with $\acute{y}_{m}^{\textbf{s}}$ from the feature embeddings $h_{m}^{\textbf{s}}$. Furthermore, we introduce a fixed Identity-head $I_{m}$ head composed of two fully connected layers, as shown in Figure 5. It takes embeddings $\mathbf{h}_{m}^{\textbf{s}}$ without gradient backpropagation. $I_{m}$ works as a regularizer in conjunction with the disentanglement loss to suppress non-discriminative identity-related information during target adaptation. The disentanglement loss decouples identity-related information, while the $I_{m}$ helps to constrain redundant (non-task-specific) features, ensuring the model focuses on learning relevant expression-related representations.
9 Datasets
9.1 Biovid Heat Pain
[walter2013biovid] dataset consists of 87 subjects captured in a controlled environment. It consists of five classes, including: No-Pain, PA-1, PA-2, PA-3, and PA-4. Every individual recorded 20 videos per class, corresponding to 100 videos per subject. We follow the same protocol as [werner2017analysis], where it eliminates the initial 2 seconds from the video, which does not show any spike indicating pain. In our experiment, we follow [zeeshan2024subject] to consider 77 subjects as sources and 10 subjects as targets: Sub-1 (081014_w_27), Sub-2 (101609_m_36), Sub-3 (112009_w_43), Sub-4 (091809_w_43), Sub-5 (071309_w_21), Sub-6 (073114_m_25), Sub-7 (080314_w_25), Sub-8 (073109_w_28), Sub-9 (100909_w_65), Sub-10 (081609_w_40). The label distribution statistics for the test set of Biovid were shown in Tab. 5. Each subject has nearly balanced distributions across the five classes (neutral + 4 pain levels), confirming the dataset is overall class-balanced.
9.2 StressID
[chaptoukaev2023stressid] dataset comprised 65 (47 men and 18 women) subjects taken in a lab-controlled environment. Each participant was exposed to 11 tasks, grouped into 4 categories: watching emotional videos, breathing, interactive tasks, and relaxation. These tasks were captured using three different modalities (visual, physiological, and audio). Due to the unavailability of certain modalities in 11 of the participants, we only consider 54 individuals, which includes all the modalities. In our experiments, we take 44 subjects as sources and 10 subjects as targets. The target subject selection follows a 70% (men) and 30% (women) ratio, reflecting the higher number of male participants in the dataset. For each participant, there are up to 11 task-specific videos, from which we extracted frames at 1 fps for per-image expression classification. Tab. 6 shows the selected target subject demographics (ID, Gender) and the total number of samples. Tab. 7 shows the label distribution of StressID datasets. It exhibits a noticeable imbalance between stress and no-stress samples across subjects. Only two subjects, Sub-1 and Sub-8, are somewhat balanced, while others are skewed.
9.3 Behavioural Ambivalence/Hesitancy (BAH)
The Behavioural Ambivalence/Hesitancy (BAH) [gonzalez2025bah] dataset is the first multimodal resource designed for subject-specific recognition of ambivalence and hesitancy (A/H) in real-world settings. It consists of 1,118 videos (8.26 hours, including 1.5 hours of A/H) from 224 participants across 9 Canadian provinces, covering diverse ages and ethnicities. Participants answered 7 questions, some explicitly eliciting A/H, while being recorded via webcam and microphone, providing multimodal signals such as facial expressions, vocal cues, and transcripts. Expert annotations mark A/H occurrences at both frame- and video-level, with aligned face crops, transcripts, and participant metadata also included. Unlike controlled lab datasets (e.g., BioVid, StressID), BAH captures natural, in-the-wild behaviours, making it uniquely challenging and valuable for multimodal domain adaptation. In our experiments, we use 143 training data from BAH as subjects to consider as source domains and use 5 subjects from the test-set of BAH as targets: Sub-1 (82711), Sub-2 (82585), Sub-3 (82683), Sub-4 (82708), and Sub-5 (82632), enabling a realistic evaluation of MuSACo under noisy, heterogeneous conditions. The label distribution statistics for the test set of Stress-ID were shown in Tab. 8 It also shows a significant imbalance in ambivalence vs. non-ambivalence across most subjects, with only Sub-1 (82711) and Sub-8 (82683) being closer to balanced.
Subjects ID NA PA1 PA2 PA3 PA4 Total
Sub-1 081014_w_27 287 285 311 305 312 1500
Sub-2 101609_m_36 325 278 323 274 300 1500
Sub-3 112009_w_43 304 282 296 310 308 1500
Sub-4 091809_w_43 265 316 295 298 302 1476
Sub-5 071309_w_21 299 297 295 304 297 1492
Sub-6 073114_m_25 295 310 320 283 292 1500
Sub-7 080314_w_25 311 318 288 276 307 1500
Sub-8 073109_w_28 314 300 300 272 303 1489
Sub-9 100909_w_65 298 298 313 279 312 1500
Sub-10 081609_w_40 300 297 313 286 303 1499
Table 5: Label distribution for BioVid dataset across target subjects test-set.
Subjects ID Gender NSV SV No. frames
Sub-1 kycf Men 5 6 1021
Sub-2 uymz Men 8 3 880
Sub-3 h8s1 Men 4 7 957
Sub-4 ctzy Men 2 9 1041
Sub-5 p9i3 Men 5 6 1041
Sub-6 7h5u Men 5 5 917
Sub-7 g7r2 Men 8 3 1041
Sub-8 b9w0 Women 4 7 1026
Sub-9 r3zm Women 3 8 909
Sub-10 x1q3 Women 2 9 1001
Table 6: StressID target subject demographics and total number of samples (NSV: No Stress Videos, SV: Stress Videos)
Subjects ID No-Stress Stress Total
Sub-1 kycf 118 86 204
Sub-2 uymz 148 28 176
Sub-3 h8s1 128 83 211
Sub-4 ctzy 60 148 208
Sub-5 p9i3 125 83 208
Sub-6 7h5u 127 57 184
Sub-7 g7r2 175 33 208
Sub-8 b9w0 105 100 205
Sub-9 r3zm 63 120 183
Sub-10 x1q3 59 142 201
Table 7: Label distribution for StressID dataset across target subjects test-set.
Subjects ID No-Amb Amb Total
Sub-1 82711 433 874 1307
Sub-2 82687 319 207 526
Sub-3 82585 575 85 660
Sub-4 82592 737 93 830
Sub-5 82598 1199 189 1388
Sub-6 82632 854 243 1097
Sub-7 82681 354 207 561
Sub-8 82683 445 456 901
Sub-9 82708 249 104 353
Sub-10 82714 231 234 465
Table 8: Label distribution for BAH dataset across target subjects test-set.
10 Ablation
10.1 Impact of Weighting Hyperparameters.
We performed a sequential weight sensitivity analysis for the key loss terms, $\mathcal{L}_{unsup}^{t}$, $\mathcal{L}_{aw}$, and $\mathcal{L}_{agn}$ shown in Tab. 9). The weight sensitivity analysis was performed on three representative subjects from the BioVid dataset: a young woman (081014_w_27), a young man (073114_m_25), and an older woman (100909_w_65). The reported results were the average of all three subjects, and each experiment was run for 10 epochs. These subjects were chosen to reflect the diversity of the dataset across age and gender groups. First, we varied the weight of the target PL loss while keeping the others fixed at 1. Using the best result, we tuned the class-aware alignment loss, followed by the class-agnostic loss, each time fixing the previously selected best weights. This sequential tuning strategy allowed us to identify the most effective configuration while minimizing computation cost. The best configuration obtained was $\gamma_{t}=1$, $\alpha=0.1$, and $\beta=0.5$. All experiments in the paper were conducted using this configuration to ensure consistency across results.
Weights Accuracy
$\mathcal{L}_{unsup}^{t}$ $\mathcal{L}_{aw}$ $\mathcal{L}_{agn}$
0.01 38.6 40.7 40.9
0.05 40.5 40.7 43.8
0.1* 40.3 43.9 41.4
0.3 39.9 39.9 40.1
0.5* 38.7 40.3 44.1
0.7 39.9 39.9 39.9
0.9 40.2 41.4 43.9
1* 40.7 40.7 40.9
2 40.6 39.6 41.9
5 40.2 38.8 22.0
Table 9: Ablation study on loss weights. Each loss weight is varied while keeping the previously selected optimal weights fixed.
$\tau_{ss}$ Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg SS Avg Acc
0.00 77 77 77 77 77 77 77 77 77 77 77.0 37.66
0.05 56 55 62 52 57 50 59 53 52 65 56.1 37.59
0.10 53 50 57 52 53 48 57 52 49 63 53.4 38.02
0.15 50 47 56 49 49 45 52 50 44 61 50.3 33.92
0.20 48 44 50 47 45 43 48 44 40 54 46.3 36.51
0.25 43 39 44 44 45 43 41 40 35 51 42.5 36.70
0.30 39 35 40 37 43 41 37 38 33 49 39.2 38.90
0.35 35 28 34 35 38 36 32 36 29 44 34.7 35.52
0.40 31 22 28 32 35 33 30 28 27 39 30.5 37.27
0.45 31 20 26 26 31 30 25 20 21 35 26.5 38.39
0.50 26 16 24 26 28 26 21 16 18 34 23.5 38.69
0.55 20 23 23 24 22 23 16 13 15 30 20.9 39.89
0.60 16 15 21 20 15 18 16 10 12 27 17.0 38.29
0.65 12 13 20 19 15 16 11 11 10 22 14.9 37.62
0.70 9 8 15 15 12 15 7 6 9 21 11.7 37.48
0.75 7 6 13 12 10 13 5 5 8 20 9.9 37.42
0.80 5 5 9 7 6 11 3 4 8 16 7.4 31.31
0.85 4 4 6 7 5 9 3 3 5 13 5.9 30.27
0.90 2 2 6 6 4 7 3 3 3 10 4.6 27.93
0.95 2 2 4 3 3 3 2 2 3 5 2.9 28.53
1.00 2 2 2 2 2 2 2 2 2 2 2.0 29.12
Table 10: Number of selected subjects ($S$) for each target subject across different $\tau_{ss}$ thresholds. Lower $\tau_{ss}$ values indicate fewer selected sources. We also report the average number of selected subjects and the corresponding average accuracy.
10.2 Source Subject Selection Analysis
Tab. 10 shows the number of selected source subjects for each of the 10 target subjects for the Biovid dataset for various threshold ($\tau_{ss}$) values. As $\tau_{ss}$ increases, fewer source subjects are selected, reducing potential noise from irrelevant domains. Notably, the accuracy improves when transitioning from $\tau=0$ to moderate thresholds (e.g., $\tau_{ss}=0.45$ or $0.55$), demonstrating that selective inclusion of source data benefits generalization. The best average accuracy of 39.89% is achieved with $\tau_{ss}=0.55$, using approximately 21 subjects on average. To assess the robustness of the source subject selection threshold ($\tau_{ss}$), we conducted experiments on the Behavioural Ambivalence/Hesitancy (BAH) [gonzalez2025bah] dataset, which differs substantially from lab-controlled datasets (BioVid and Stress) as it consists of self-recorded, in-the-wild videos. Table 11 shows results across thresholds from 0.45 to 0.65. The analysis reveals that accuracy remains stable in the 0.45–0.55 range, with the best performance achieved at $\tau_{ss}=0.55$. Higher thresholds reduce the number of selected sources and lead to performance degradation due to over-pruning. These findings confirm that $\tau_{ss}$ is not highly sensitive to fine-tuning and that a default value around 0.55 provides a reliable trade-off between accuracy and source diversity across datasets.
10.3 Ablation on Target PLs Threshold ($\tau_{pl}$).
Tab. 12, and found that $\tau_{0}=0.95$ provided the best trade-off between label quality and sample quantity. Rather than fixing this value, we employ a confidence-annealing schedule, where $\tau_{pl}$ is decreased by $0.01$ every $N$ epochs, allowing more target samples to be included as the model stabilizes. This ensures early training benefits from high-confidence labels, while later stages incorporate a broader set of samples. To further mitigate the exclusion of useful but uncertain samples, we introduce a domain-agnostic loss that allows learning from low-confidence examples as well. Target ground-truth labels were only used here to evaluate pseudo-label quality.
$\tau_{ss}$ Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg SS Avg Acc
0.45 44 33 39 41 39 39.2 67.1
0.50 42 31 36 40 36 37.0 68.2
0.55 41 30 35 38 34 35.6 68.9
0.60 39 28 33 37 33 34.0 66.6
0.65 38 27 30 36 30 32.2 62.5
Table 11: Impact of varying source subject selection threshold $\tau_{ss}$ on the BAH dataset.
Threshold ($\tau_{pl}$) Accuracy
0.50 40.2
0.60 48.1
0.70 44.7
0.75 46.5
0.80 46.2
0.85 48.6
0.90 46.7
0.95 52.0
Table 12: Accuracy at different pseudo-label confidence thresholds ($\tau_{pl}$)
Exp-head Id-head Knife Visual Physio
25.1 36.2
26.3 37.0
28.3 38.2
Table 13: Ablation on the disentanglement module for visual and physiological modalities for the Biovid dataset.
10.4 Impact of Disentanglement
conducted separately for visual and physiological modalities on the BioVid dataset, using 77 source and 10 target subjects. Results (Fig. 6, Tab. 13) show that incorporating the KNIFE-based disentanglement improves performance for both modalities. Without disentanglement, identity information interferes with expression classification, resulting in reduced accuracy. Adding the KNIFE-loss mitigates this, and introducing an identity head yields further improvements, confirming that explicitly removing identity bias enhances modality-specific expression features.
10.5 Computational Complexity
[…]
Figure 6: Impact of adding disentanglement.
Method Src Train Trg Train Inf Acc
CAN [kang2020contrastive] 59 19,625 8.03 34.6
Sub-based [zeeshan2024subject] 59 234 7.36 34.6
CMSDA [scalbert2021multi] 271 4.07 36.8
Ours (all srcs) 59 1895 7.82 39.3
Ours (co-train) 59 426 7.64 43.8
Table 14: Training and inference time (per batch) of our proposed and baseline methods. All times are shown in ms.
$\mathcal{L}^{s}$ $\mathcal{L}^{t}$ $\mathcal{L}_{aw}$ $\mathcal{L}_{agn}$ Avg Acc
32.1
36.5
38.0
37.2
40.7
Table 15: Impact of individual loss components under equally weighted hyperparameter.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
63.1
36.6
57.3
24.6
45.6
32.1
61.2
37.7
50.0
37.8
22.3
36.6
58.3
37.5
51.5
50.8
40.8
63.4
43.2
45.6
68.4
36.6
33.8
39.5
46.9
25.6
52.6
54.8
22.7
53.4
47.7
34.8
50.5
$M$-UDA
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
61.2
57.1
57.1
37.6
45.3
43.2
54.0
49.3
55.6
64.3
61.3
74.1
51.3
46.6
40.3
86.2
53.6
48.3
45.6
55.7
56.6
58.3
55.6
61.5
56.6
59.6
55.6
64.8
47.6
45.6
55.6
53.4
62.3
62.2
55.4
74.4
38.6
39.6
49.3
50.0
52.9
52.5
53.0
60.7
$M$-MSDA
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
65.3
68.9
76.1
45.4
45.3
59.6
54.6
74.2
69.3
42.3
46.6
53.3
51.2
55.2
58.6
58.6
62.6
67.2
66.6
56.6
75.2
54.2
39.3
48.3
59.3
52.2
74.6
34.6
39.9
40.2
53.2
54.1
62.2
Upper Bound Fine-tuning 86.3 43.8 94.4 97.7 79.6 75.6 45.6 93.0 91.0 97.0 80.4
Table 16: Subject-wise F1 scores of StressID dataset across 10 target subjects.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
29.8
38.1
41.2
26.1
32.5
27.2
9.1
39.1
24.9
28.1
15.4
32.2
19.3
10.4
29.5
21.6
29.0
25.8
25.7
14.3
34.7
15.3
34.6
24.3
16.5
28.7
22.3
16.1
22.0
18.9
20.8
26.4
28.1
$M$-UDA
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
32.5
28.5
28.3
35.6
25.6
24.3
20.7
21.7
29.5
25.5
29.6
33.3
33.6
29.2
32.4
36.2
28.3
25.2
28.4
28.7
25.1
24.2
29.5
16.0
29.6
28.6
32.1
31.2
24.2
23.4
28.6
29.4
25.2
23.1
25.4
35.4
20.2
20.5
26.5
17.6
27.4
25.3
28.2
28.5
$M$-MSDA
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
CMSDA [scalbert2021multi]
MuSACo (MSDA)
34.6
34.6
31.2
41.6
30.3
37.6
32.5
33.2
22.3
22.4
21.2
49.6
24.6
20.5
31.2
27.8
27.9
34.5
32.3
25.6
26.3
24.6
30.1
36.5
28.6
36.8
30.2
29.4
26.6
24.3
34.1
35.4
33.2
30.6
38.9
49.0
18.1
17.6
18.5
29.2
27.3
28.4
30.0
35.7
Upper Bound Fine-tuning 81.2 74.3 67.6 68.2 70.3 61.4 74.2 70.6 74.0 49.5 69.1
Table 17: Subject-wise F1 scores of BioVid dataset on 10 target subjects.
Settings Methods Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg F1
Lower Bound Visual-only 25.2 46.2 35.2 41.3 43.2 38.2
Audio-only 27.4 34.4 34.1 40.3 44.1 36.1
Fusion 26.3 44.5 34.7 41.9 44.6 38.4
$M$-UDA MMD [sejdinovic2013equivalence] 28.2 44.5 39.3 38.7 43.2 38.8
MuSACo (UDA) 26.1 46.1 35.6 42.2 43.5 38.7
$M$-MSDA Sub-basedtop-k [zeeshan2024subject] 24.6 45.4 36.8 42.7 44.9 38.9
MuSACo (MSDA) 28.2 48.6 39.4 44.2 45.4 41.2
Table 18: Subject-wise F1 scores on the BAH dataset across five target subjects. MuSACo MSDA consistently outperforms baselines and other UDA/MSDA methods.
10.6 Impact of Difference Loss Components with Equal Weights
In addition to the tuned-weight ablation reported in the main paper, we also conducted a diagnostic experiment where all loss weights were fixed to 1. This isolates the relative impact of each component without any hyperparameter tuning. Table 15 shows that the full model under equal weights reached 40.7%, which is lower than the tuned configuration (43.8%). Importantly, the incremental trend remains consistent across both setups: $\mathcal{L}^{t}$, $\mathcal{L}_{aw}$, and $\mathcal{L}_{agn}$ each contribute positively to performance. This confirms that the effectiveness of MuSACo does not hinge on carefully tuned weights, but rather that appropriate weighting further amplifies the gains. Together, these results underline the robustness and general applicability of the proposed loss formulation.
10.7 Target Pseudo-label Progression During Training
We evaluate the reliability of co-training by tracking pseudo-label (PL) accuracy during training. Results are reported for two challenging StressID subjects: ctzy (Sub-4) and x1q3 (Sub-10), which exhibit high label imbalance. To stabilize PL quality, we adopt a confidence-annealing schedule, initializing $\tau_{pl}=0.95$ and gradually reducing it by 0.01 every $N$ epochs (e.g., 0.94 at Ep-4, 0.93 at Ep-8). As shown in Fig. 7, PL accuracy steadily improves across epochs, demonstrating the effectiveness of iterative refinement under this schedule. For Sub-4, PL accuracy improves from 0.976 at initialization to 0.996 after 8 epochs, while for Sub-10, which is highly imbalanced, accuracy increases from 0.602 to 0.666. These trends indicate that co-training with confidence annealing reliably enhances PL quality, even under challenging conditions.
Figure 7: Target pseudo-labels accuracy during training.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg Acc
Lower Bound
Visual-only
Physio-only
Fusion
37.8
41.7
36.0
27.7
30.1
28.0
23.3
30.5
26.0
22.6
33.3
35.0
27.5
36.7
35.0
21.6
30.9
15.0
37.0
34.7
43.0
20.0
38.0
40.0
22.0
37.3
35.0
24.0
28.0
28.0
26.4
34.1
32.1
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
40.3
39.0
31.2
35.0
32.3
34.0
26.3
32.0
33.3
29.0
24.1
19.0
34.3
38.9
39.5
37.0
37.3
31.0
20.0
29.9
31.9
32.5
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
45.0
42.0
35.0
38.0
39.0
51.0
37.0
30.0
34.0
28.0
32.0
39.0
42.0
43.0
36.0
35.0
27.0
43.0
30.0
33.0
35.7
38.2
Table 19: Subject-wise accuracy of the BioVid dataset using the ViT backbone across 10 target subjects.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
35.6
31.0
23.9
27.5
21.9
27.1
22.5
23.1
22.8
12.5
24.5
22.6
19.7
27.9
33.5
19.6
23.6
9.5
35.3
23.9
32.9
15.6
27.3
34.6
14.9
28.2
29.1
23.2
19.1
26.1
22.6
25.0
26.2
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
31.7
33.6
33.1
34.6
28.6
25.9
20.5
26.9
33.8
25.9
23.5
18.4
37.7
38.4
35.6
33.8
29.3
31.3
18.1
29.8
29.2
29.9
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
36.9
33.9
27.2
27.6
34.6
38.7
32.9
26.4
20.6
24.9
27.4
33.9
34.5
41.5
28.0
30.1
13.1
37.6
22.6
29.8
27.8
32.4
Table 20: Subject-wise F1 scores of the BioVid dataset using the ViT backbone across 10 target subjects.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg Acc
Lower Bound
Visual-only
Physio-only
Fusion
57.0
57.8
48.3
18.3
84.1
14.0
41.7
60.7
49.3
68.3
28.8
78.3
62.4
60.1
66.0
47.5
69.0
49.2
84.2
84.1
74.0
47.0
51.2
51.3
65.5
34.4
64.2
69.7
29.4
73.5
56.2
56.0
56.8
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
53.0
64.0
23.0
20.0
36.4
38.0
82.0
79.0
69.0
65.0
47.6
50.0
78.0
82.0
46.6
50.0
77.0
66.0
66.0
68.0
57.9
58.2
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
58.0
38.0
15.0
38.0
50.0
87.0
72.0
100.0
62.0
67.0
82.0
67.0
94.0
85.0
82.0
82.0
66.0
66.0
67.0
68.0
64.8
69.8
Table 21: Subject-wise accuracy of the StressID dataset across 10 target subjects using the ViT backbone.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
55.2
36.6
47.0
17.3
45.7
13.0
32.5
37.8
36.4
67.9
22.4
77.4
46.1
37.5
53.4
46.3
40.8
47.1
76.4
45.7
64.2
32.9
33.9
42.8
39.6
25.6
39.6
41.1
22.7
42.3
45.5
34.9
46.3
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
51.6
62.6
22.6
19.3
27.5
28.5
72.6
77.4
55.2
49.3
42.1
48.5
72.5
72.8
44.6
34.8
44.0
39.5
39.5
40.2
47.2
47.3
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
56.0
28.5
13.8
36.4
46.5
85.2
71.0
98.8
38.0
66.3
80.1
66.2
87.3
45.7
81.2
81.2
39.5
39.5
39.9
40.2
55.3
58.8
Table 22: Subject-wise F1 scores of the StressID dataset across 10 target subjects using the ViT backbone.
11 Subject-wise F1 Scores
11.1 Stress-ID
Table 16 presents subject-wise F1 scores on the StressID dataset. While the dataset has a binary label distribution (stress vs. no-stress), clear variability emerges across subjects. For example, lower-bound fusion consistently yields higher F1 than either modality alone, showing the complementary nature of visual and physiological signals. However, without adaptation, predictions remain unbalanced across subjects, with several experiencing sharp drops (e.g., Sub-2, Sub-4). UDA methods improve F1 moderately on some subjects, but results remain inconsistent. In contrast, MuSACo (MSDA) achieves the best overall performance, boosting recall on harder subjects (e.g., Sub-9) and producing the highest average F1 across all targets. These findings underscore the importance of F1 as a complementary metric to accuracy, particularly for stress recognition, where subject-specific imbalance strongly affects model robustness.
11.2 Biovid
Tab. 17 reports subject-wise F1 scores on the BioVid dataset. Our analysis shows that MuSACo particularly struggles on subjects where the Physio-only baseline itself is weak (e.g., Sub-4, Sub-5, Sub-7). This suggests that physiological variability across individuals is the key limiting factor: when physio cues are unreliable, visual signals alone are too subtle to compensate, and fusion inherits this weakness. These subject-specific outcomes highlight an inherent challenge of BioVid, where the modality that is usually strongest (physio) can be inconsistent for some individuals, constraining overall adaptation performance. Nevertheless, MuSACo achieves consistent performance across all subjects and outperforms competing methods on 6 out of 10 targets, leading to the best overall F1 score.
11.3 BAH
Tab. 18 reports subject-wise F1 scores on the BAH dataset. While the visual modality is generally stronger than audio, naive fusion provides slight gains by leveraging complementary information across modalities. MuSACo (MSDA) achieves the best overall performance, consistently surpassing both lower-bound baselines and existing UDA/MSDA approaches, demonstrating the effectiveness of source subject selection and multimodal adaptation in more challenging, in-the-wild conditions.
12 Evaluation on Additional Backbone
In the main paper, all experiments are conducted using a ResNet-18 backbone to ensure fair comparison with prior work and maintain a computationally efficient setup. However, MuSACo is inherently backbone-agnostic and does not rely on architectural properties specific to convolutional networks. To verify that MuSACo extends beyond ResNet-based architectures, we additionally evaluate the method using a Vision Transformer (ViT). We use the same MuSACo protocol as before for all the experiments, including the same setup for baseline methods. Biovid. Tables 19 and 20 report the subject-wise Accuracy and F1 scores across 10 target subjects. MuSACo consistently surpasses the baseline UDA and MSDA methods, achieving the highest average Accuracy and F1. Notably, MuSACo maintains strong performance even in subjects where the baselines degrade, demonstrating stable target adaptation. StressID. Tables 21 and 22 present the subject-wise Accuracy and F1 results. MuSACo consistently outperforms the baseline UDA and MSDA methods across most subjects, achieving the highest average Accuracy and F1. In particular, MuSACo demonstrates strong improvements on challenging subjects such as Sub-3, Sub-4, and Sub-6, where existing approaches show significant performance drops. These results further validate that the effectiveness of MuSACo is preserved even when replacing ResNet-18 with a transformer-based backbone, confirming that the proposed method is robust and generalizes well across different architectural families.
13 MuSACo
MuSACo (inspired by the layered harmony of Musaca) integrates two synergistic modules for subject-specific adaptation: a co-training-based source subject selection module that identifies the most relevant sources using complementary cues from multiple modalities, and an adaptation module that aligns source and target domains using class-aware and class-agnostic losses. Like Musaca’s layered composition—where distinct elements work in harmony, MuSACo leverages modality-specific strengths to guide pseudo-label generation and fuse information effectively. This enables robust, personalized adaptation for each target subject while maintaining consistency across modalities.